# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ 3.3. Learning algorithms
[21] If a function $f$ is solvable, satisfies the strong ascending chain condition and has a progress measure, then the learning algorithm from Figure LABEL:fig:algo terminates and returns an automaton computing $f$.
4. Hierarchy of learnable functions
Figure 6. General hierarchy.
In the previous section, we have defined the classes of strongly guessable, guessable, and weakly guessable functions, and in Appendix A the dual notions of strongly co-guessable, co-guessable, and weakly co-guessable functions. We obtain a general hierarchy depicted in Figure 6. Next we show that parts of this hierarchy collapse in some natural settings, such as for commutative semirings (see Figure 7 below) and give examples to demonstrate that certain of these containments can be strict. The examples are drawn from various well-studied semirings discussed in Section 2, but one could obtain a unified semiring separating these classes by considering their product with component-wise operations.
4.1. Commutative semirings
For commutative semirings we have the following result, giving the hierarchy depicted in Figure 7.
Figure 7. Hierarchy for commutative semirings
@@ 4.1. Commutative semirings
Let $S$ be a commutative semiring and let $f:\Sigma^{*}\rightarrow S$ be computed by a weighted automaton over $S$. If $f$ is weakly guessable and weakly co-guessable, then $f$ is strongly guessable and strongly co-guessable.
@@
Suppose that $f$ is weakly guessable and weakly co-guessable and let $F$ be the Hankel matrix of $f$. We first show that $f$ is strongly guessable. Since $f$ is weakly co-guessable, we may fix a finite set of words $T$ such that the columns indexed by $T$ are column-generating. We will prove that for any finite set of words $Q$ with $\Lambda_{Q}\neq\emptyset$ we have $\Lambda_{Q}=\Lambda_{Q,T}$.
Thus let $Q$ be arbitrary with the property that $\Lambda_{Q}\neq\emptyset$ and let $\lambda\in\Lambda_{Q,T}$. Then for all $q,p\in Q$, all $a\in\Sigma$ and all $t\in T$ we have: $F_{qa,t}=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes F_{p,t}$. By our assumption on $T$ we know that for all words $w$ and all $t\in T$ there exist $\mu_{w,t}\in S$ such that $F_{u,w}=\bigoplus_{t\in T}\mu_{w,t}\otimes F_{u,t}$ holds for all words $u$. In particular, for all $w\in\Sigma^{*}$, we have:
$F_{qa,w}$ $=$ $\bigoplus_{t\in T}\mu_{w,t}\otimes F_{qa,t}=\bigoplus_{t\in T}\mu_{w,t}\otimes(\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes F_{p,t})$
$=$ $\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes(\bigoplus_{t\in T}\mu_{w,t}\otimes F_{p,t})$
$=$ $\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes F_{p,w}.$
In other words, $\langle qa\rangle=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes\langle p\rangle$, giving $\lambda\in\Lambda_{Q}$, as required. This demonstrates that $f$ is strongly guessable.
Since $f$ is weakly guessable and co-guessable, then $\bar{f}$ is also weakly guessable and weakly co-guessable. The above argument therefore yields that $\bar{f}$ is strongly guessable and so $f$ is strongly co-guessable. ∎
We conclude this section with the following useful observations concerning commutative semirings:
@@
We next aim to show that for commutative semirings almost all inclusions depicted in Figure 7 can be strict.
4.2. Examples for the strict hierarchy.
The full proofs of the membership of the following functions to certain parts of the hierarchy are given in Appendix B. We give here a quick overview.
For some semirings $S$, there are functions $f:\Sigma^{*}\rightarrow S$ that are neither weakly guessable, nor weakly co-guessable. We show that this is the case for $f_{1}$ computed by the automaton given in Figure 1(b), whether viewed over $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$, $f^{\prime}_{1}$ computed by the automaton given in Figure 1(d), whether viewed over $\mathbb{N}$ or $\mathbb{R}_{\geq 0}$, and $f^{\prime\prime}_{1}$ computed by the automaton in Figure 1(f). There are also functions that are weakly guessable but neither guessable nor weakly co-guessable. We show that this is the case for $f_{2}$ computed by the automaton given in Figure 1(a), whether viewed as an automaton over $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$. It then follows that the mirror function $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable.). Finally, we show that the functions $f_{3}$ computed by the automaton depicted in Figure 1(a) (over $\mathbb{N}_{\max}$), $f^{\prime}_{3}$ computed by the automaton from Figure 1(c) and $f^{\prime\prime}_{3}$ computed by the automaton from Figure 1(e) are all strongly guessable but not weakly co-guessable. The functions $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable. It is in general easy to find examples at the intersection of all the classes: for commutative semirings, any function computed by an automaton that is both literal and co-literal lies in this intersection. A trivial example of such function would be the constant function $f_{4}$ mapping all words to $1_{S}$.
@@ 4.3. Collapse in some familiar semirings
In many well-studied semirings, we note that the hierarchy collapses further, due to the fact that certain ascending and descending chain conditions are satisfied. It is clear from the results of previous section that the following two properties characterise (A) when all functions computed by a finite state weighted automaton over $S$ are weakly guessable and weakly co-guessable (as in Figure 8(a)) and (B) when all weakly guessable functions are strongly guessable and all weakly co-guessable functions are strongly co-guessable (as in Figure 8(b)):
@@
Clearly if both the conditions depicted in Figures 8(a) and 8(b) hold, we have that all functions computed by finite weighted automata are strongly guessable. Note that this is the case in the semirings $\mathbb{Z}$, $\mathbb{R}$ and $\mathbb{B}$ (as special cases of principal ideal domains, fields and finite semiring respectively) for example. Further, the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$ and $\mathcal{P}_{\rm fin}(\Sigma^{*})$ satisfy Property B:
@@
We give the proof for rows; he proof for columns is dual. Let $S$ be one of the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$, or $\mathcal{P}_{\rm fin}(\Sigma^{*})$. Let $f$ be a weakly guessable function computed over $S$, and $F$ its Hankel matrix. Let $Q$ be such that $\Lambda_{Q}$ is non-empty. First, note that, for some fixed $T$, if no row indexed by $Q$ and restricted to $T$ is fully $0_{S}$, then in the three cases, $\Lambda_{Q,T}$ is finite. Let $Z_{Q,T}=\{q\in Q\mid\langle q\rangle_{T}\mbox{ is a }0_{S}\mbox{ row}\}$, $Q^{\prime}=Q\setminus Z_{Q,T}$ and $N_{Q^{\prime},T}$ the set of tuples $(\lambda_{p}^{\varepsilon},\lambda_{q,a,p})_{q\in Q,a\in\Sigma,p\in Q^{\prime}}$ from $\Lambda_{Q,T}$ where $p$ is restricted to $Q^{\prime}$. Clearly $Z_{Q,T}$ is finite and its size is bounded above by the one of $Q$. Notice that $N_{Q^{\prime},T}$ is also a finite set, since each $\lambda_{q,a,p}$ where $p\in Q^{\prime}$ can only take a finite number of different values. Suppose that $T\subsetneq T^{\prime}$ with $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$. It is clear from the definition that $Z_{Q,T}\supseteq Z_{Q,T^{\prime}}$, and either this containment is strict, or if not, then $N_{Q^{\prime},T}\supsetneq N_{Q^{\prime},T^{\prime}}$. Since both sets have a finite size, there cannot be an infinite descending chain as in the definition. ∎
@@
Suppose that $f:\Sigma^{*}\rightarrow S^{\prime}$ is weakly guessable. Thus there is a literal automaton $\mathcal{A}$ with weights in $S^{\prime}$ that computes $f$. Viewed as a function over $S$, we also have that $f$ is weakly guessable (since $\mathcal{A}$ is a literal automaton with weights in $S$ computing $f$) and hence by assumption, strongly guessable.
Suppose for contradiction that $f$ not strongly guessable over $S^{\prime}$. Then there exists a finite set $Q$ such that $\Lambda_{Q}\neq\emptyset$ and for all finite sets $T$ there exists $\lambda_{T}\in\Lambda_{Q,T}\setminus\Lambda_{Q}$. However, noting that the coefficients $\lambda_{T}$ lie in $S$, this directly contradicts that $f$ is strongly guessable over $S$.
The final statement follows from Remark 3.9 together with the fact that any integral domain embeds into a field (its field of fractions). ∎
The table in Figure 9 demonstrates that there are interesting and important semirings (including the natural numbers) where the properties of weakly guessable and strongly guessable coincide (i.e. Property B holds) whilst there exist functions that are not weakly guessable (i.e. Property A does not hold). Since the property of being weakly guessable is characterised in terms of automata by Theorem 3.5, over such semirings it is possible for the ‘teacher’ to select a guessable function to begin with. Finally, we say that semirings in which all six classes of guessability collapse have property $D$. Clearly if Properties A and B both hold, then D must also hold. Conversely, if Property B does not hold, then Property D cannot hold. We do not have an example where Propoerty D holds but Property A does not.
A B D
$\mathbb{B}$ Yes Yes Yes
$\mathbb{R}$ Yes Yes Yes
$\mathbb{Z}$ Yes Yes Yes
$\mathbb{N}$ No ($f^{\prime}_{1}$) Yes No ($f^{\prime}_{3}$)
$\mathbb{R}_{\geq 0}$ No ($f^{\prime}_{1}$) Yes No ($f_{5}$)
$\mathcal{P}_{\rm fin}(\Sigma^{*})$ No ($f^{\prime\prime}_{1}$) Yes No ($f^{\prime\prime}_{3}$)
$\mathbb{N}_{\max}$ No ($f_{1}$) Yes No ($f_{3}$)
$\mathbb{Z}_{\max}$ No ($f_{1}$) No ($f_{2}$) No
$\mathbb{R}_{\max}$ No ($f_{1}$) No ($f_{2}$) No

### v4

@@ 3.3. Algorithmic strategies
[36, Theorem 14] Suppose that a closure strategy exists for $f$. If $f$ satisfies the ascending chain condition on rows and has a weak progress measure, then the learning algorithm from Figure LABEL:fig:algo terminates and returns an automaton computing $f$.
Remark 3.26.
Finally, we note the relationship between our notions of guessability and the termination conditions stated above.
  • The ascending chain condition on rows is weaker than the ascending chain condition considered in [36] (which considers all chains of submodules, including those which may not be generated by finitely many rows of the matrix) but stronger than the weak ascending chain condition from Definition 3.4 (which considers only certain chains of submodules generated by finitely many rows of the matrix). Thus a function satisfying the termination conditions above is certainly (by Theorem 3.6) weakly guessable. Weak guessability alone is not sufficient to ensure termination of algorithm from Figure LABEL:fig:algo (as we have seen, it is precisely the condition required in order for a correct hypothesis automaton to exist).
  • The existence of a progress measure (in the sense of [36]) for a Hankel matrix implies that the corresponding function is strongly row-bound. However, the converse need not be true. A progress measure implies a uniform bound, for all $Q$, on the number of elements in a total $Q$-chain, but for a (strongly) row-bound function, even if no such chain can be infinite, its size might depend on the set of words $Q$.
  • Each progress measure gives rise to a weak progress measure, and moreover, if a function has a weak progress measure then it is strongly row-bound or equivalently (by Proposition 3.13) strongly guessable. Thus the proposition above guarantees that the learning algorithm in Figure LABEL:fig:algo terminates for a class of strongly guessable functions. Strong guessability alone is not sufficient to ensure termination of algorithm from Figure LABEL:fig:algo. However, it follows from the proof of Proposition 3.19 that there is an algorithm with a different (incremental) update strategy for the sets $Q$ and $T$ that will terminate with success on any strongly guessable function for which a closure strategy exists.
4. Hierarchy of guessable functions
In the previous section, we have defined the classes of strongly guessable, guessable, and weakly guessable functions, and in Appendix A the dual notions of strongly co-guessable, co-guessable, and weakly co-guessable functions.
4.1. The general hierarchy
First, we show that for any semiring: if a function is both weakly and co-weakly guessable then it is also both strongly and co-strongly guessable, so that the general hierarchy is as depicted in Figure 3.
Figure 3. A general hierarchy for levels of guessability
@@ 4.1. The general hierarchy
Let $S$ be a semiring and let $f:\Sigma^{*}\rightarrow S$ be computed by a weighted automaton over $S$. If $f$ is weakly guessable and weakly co-guessable, then $f$ is strongly guessable and strongly co-guessable.
@@
Suppose that $f$ is weakly guessable and weakly co-guessable and let $F$ be the Hankel matrix of $f$. We first show that $f$ is strongly guessable. Since $f$ is weakly co-guessable, we may fix a finite set of words $T$ such that the columns indexed by $T$ are column-generating. We will prove that for any finite set of words $Q$, we have $\Lambda_{Q}=\Lambda_{Q,T}$.
Thus let $Q$ be arbitrary. Since $\Lambda_{Q}\subseteq\Lambda_{Q,T}$ is always true, if $\Lambda_{Q,T}=\emptyset$ then the equality holds. Otherwise, let $\underline{\lambda}\in\Lambda_{Q,T}$. Then for all $q,p\in Q$, all $a\in\Sigma$ and all $t\in T$ we have: $F_{qa,t}=\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes F_{p,t}$. By our assumption on $T$ (recalling that we consider right linear combinations in the definition of weakly co-guessable) we know that for all words $w$ and all $t\in T$ there exist $\mu_{w,t}\in S$ such that $F_{u,w}=\bigoplus_{t\in T}F_{u,t}\otimes\mu_{w,t}$ holds for all words $u$. In particular, for all $w\in\Sigma^{*}$, we have:
$F_{qa,w}$ $=$ $\bigoplus_{t\in T}F_{qa,t}\otimes\mu_{w,t}=\bigoplus_{t\in T}(\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes F_{p,t})\otimes\mu_{w,t}$
$=$ $\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes(\bigoplus_{t\in T}F_{p,t}\otimes\mu_{w,t})$
$=$ $\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes F_{p,w}.$
In other words, $\langle qa\rangle=\bigoplus_{p\in Q}\underline{\lambda}_{q,a,p}\otimes\langle p\rangle$. A similar argument shows that $\langle\varepsilon\rangle=\bigoplus_{q\in Q}\underline{\lambda}_{q}\otimes\langle q\rangle$, hence giving $\underline{\lambda}\in\Lambda_{Q}$, as required. This demonstrates that $f$ is strongly guessable. A dual argument (interchanging the rows/columns, and left/right combinations) demonstrates that $f$ is strongly co-guessable. ∎
Remark 4.2.
The proof of Lemma 4.1 actually shows that any function that is both weakly guessable and weakly co-guessable has the following stronger property. Say that a guessable function $f:\Sigma^{*}\rightarrow S$ is uniformly strongly guessable if there exists a finite set $T\subseteq\Sigma^{*}$ such that for all finite sets $Q\subseteq\Sigma^{*}$ one has $\Lambda_{Q,T}=\Lambda_{Q}$. Note the inversion of quantifiers as compared to the definition of strongly guessable; here the set $T$ has to be the same for all the sets $Q$. Dually one can define uniformly strongly co-guessable. The previous proof shows that the intersection of the class of weakly guessable and weakly co-guessable functions coincides with the intersection of the class of uniformly strongly guessable and uniformly strongly co-guessable functions. We leave open the question of whether this intersection is precisely characterised by uniform strong guessability. However, we note that it can be shown that the class of uniformly strongly guessable functions is strictly contained in the class of strongly guessable functions (as an example, the function $f_{3}$ discussed below is not uniformly strongly guessable).
4.2. Examples for the strict hierarchy.
We now give examples to demonstrate that certain inclusions in the hierarchy can be strict. The examples are drawn from various well-studied semirings discussed in Section 2, but one could obtain a unified semiring separating these classes by considering their product with component-wise operations. The full proofs of the membership of the following functions to certain parts of the hierarchy are given in Appendix B. We give here a quick overview making frequent use of the following lemma:
@@
Proof.
Let $f:\Sigma^{*}\rightarrow S$, and consider its mirror $\overline{f}:\Sigma^{*}\rightarrow S$. Let us write $F$ to denote the Hankel matrix of $f$ and $\overline{F}$ to denote the Hankel matrix of $\overline{f}$, and for a set of words $W\subseteq\Sigma^{*}$ let us also write $\overline{W}=\{\overline{w}:w\in W\}$. By definition, for all words $u,v\in\Sigma^{*}$ we have
$\overline{F}_{u,v}=\overline{f}(uv)=f(\overline{uv})=f(\overline{v}\;\overline{u})=F_{\overline{v},\overline{u}}.$
From this one sees that for any pair of finite subsets $Q,T\subseteq\Sigma^{*}$, the equations given in Notation 2.6 for the function $f$ (encoding certain left-linear combinations of the rows of $F$ indexed by $Q$ restricted to columns in $T$) are, due to commutativity of multiplication in $S$, equivalent to the equations encoding (right) linear combinations of the columns of $\overline{F}$ indexed by $\overline{Q}$ restricted to rows in $\overline{T}$, as obtained from A.1. Since weakly guessable, guessable and strongly guessable and their duals are described in terms of properties of these solution sets (considered for all finite sets $Q,T$), the result follows. ∎
For some semirings $S$, there are functions $f:\Sigma^{*}\rightarrow S$ that are neither weakly guessable, nor weakly co-guessable. We show that this is the case for $f_{1}$ computed by the automaton given in Figure 1(b), whether viewed over $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$, $f^{\prime}_{1}$ computed by the automaton given in Figure 1(d), whether viewed over $\mathbb{N}$ or $\mathbb{R}_{\geq 0}$, and $f^{\prime\prime}_{1}$ computed by the automaton in Figure 1(f). There are also functions that are weakly guessable but neither guessable nor weakly co-guessable. We show that this is the case for $f_{2}$ computed by the automaton given in Figure 1(a), whether viewed as an automaton over $\mathbb{Z}_{\max}$ or $\mathbb{R}_{\max}$. It then follows that the mirror function $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (by Lemma 4.3). Finally, we show that the functions $f_{3}$ computed by the automaton depicted in Figure 1(a) (over $\mathbb{N}_{\max}$), $f^{\prime}_{3}$ computed by the automaton from Figure 1(c) and $f^{\prime\prime}_{3}$ computed by the automaton from Figure 1(e) are all strongly guessable but not weakly co-guessable. The functions $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable. It is in general easy to find examples at the intersection of all the classes: for commutative semirings, any function computed by an automaton that is both literal and co-literal lies in this intersection. A trivial example of such function would be the constant function $f_{4}$ mapping all words to $1_{S}$.
@@ 4.3. Collapse in some familiar semirings
In many well-studied semirings, we note that the hierarchy collapses further, due to the fact that certain ascending and descending chain conditions are satisfied. It is clear from Theorem 3.6, Theorem 3.9, Proposition 3.13, Proposition 3.15 and Lemma 4.3 that the following two properties characterise (A) when all functions computed by a finite state weighted automaton over $S$ are strongly guessable and strongly co-guessable (i.e. everything collapses, as in Figure 4(a)) and (B) when all weakly guessable functions are strongly guessable and all weakly co-guessable functions are strongly co-guessable (as in Figure 4(b)):
@@
Note that Property A holds in the semirings $\mathbb{Z}$, $\mathbb{R}$ and $\mathbb{B}$ (as special cases of principal ideal domains, fields and finite semirings respectively) for example. Further, the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$ and $\mathcal{P}_{\rm fin}(\Sigma^{*})$ satisfy Property B:
@@
We give the proof for rows; the proof for columns is dual. Let $S$ be one of the semirings $\mathbb{N}_{\max}$, $\mathbb{N}$, or $\mathcal{P}_{\rm fin}(\Sigma^{*})$. Let $f$ be a weakly guessable function computed over $S$, and $F$ its Hankel matrix. Let $Q\subseteq\Sigma^{*}$ be a finite set. We aim to show that there is no total $Q$-chain. First, note that, for some fixed $T$, if no row indexed by $Q$ and restricted to $T$ is fully $0_{S}$, then in each of three semirings considered, $\Lambda_{Q,T}$ is finite (possibly empty) from which it immediately follows that there is no total $Q$-chain involving $\Lambda_{Q,T}$. Let $Z_{T}=\{q\in Q\mid\langle q\rangle_{T}\mbox{ is a }0_{S}\mbox{ row}\}$, $Q^{\prime}=Q\setminus Z_{T}$ and $N_{T}\subseteq S^{Q^{\prime}\cup(Q\times\Sigma\times Q^{\prime})}$ be the set obtained from $\Lambda_{Q,T}\subseteq S^{Q\cup(Q\times\Sigma\times Q)}$ by restricting the entries $\underline{\lambda}_{p}$ and $\underline{\lambda}_{q,a,p}$ to only permit $p\in Q^{\prime}$. Clearly $Z_{T}$ is finite and its size is bounded above by $|Q|$. Notice that $N_{T}$ is also a finite set, since each $\underline{\lambda}_{p},\underline{\lambda}_{q,a,p}$ where $p\in Q^{\prime}$ can only take a finite number of different values. Moreover, it is clear from the definitions that $T\subseteq T^{\prime}$ forces $Z_{T}\supseteq Z_{T^{\prime}}$ (i.e. any row of the Hankel matrix that is zero on all columns indexed by $T^{\prime}$ is also zero on all columns indexed by $T$), $\Lambda_{Q,T}\supseteq\Lambda_{Q,T^{\prime}}$ (by Remark 2.8) and $N_{T}\supseteq N_{T^{\prime}}$ (since these sets are obtained from the previous by restricting attention to coefficients of rows indexed by $Q$ that are not identically zero on $T$ and $T^{\prime}$ respectively). Moreover, notice that if $Z_{T}=Z_{T^{\prime}}$ (the same subset of elements of $Q$ are zero when restricted to $T$ or $T^{\prime}$) and $N_{T}=N_{T^{\prime}}$ then we must have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}$. Now, suppose that $T\subsetneq T^{\prime}$ with $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$. It is clear from the above that $Z_{Q,T}\supseteq Z_{Q,T^{\prime}}$, and either this containment is strict, or if not, then $N_{Q^{\prime},T}\supsetneq N_{Q^{\prime},T^{\prime}}$. Since both sets have a finite size, there cannot be an infinite descending chain as in the definition. ∎
@@
Suppose that $f:\Sigma^{*}\rightarrow S^{\prime}$ is weakly guessable. Thus there is a literal automaton $\mathcal{A}$ with weights in $S^{\prime}$ that computes $f$. Viewed as a function over $S$, we also have that $f$ is weakly guessable (since $\mathcal{A}$ is a literal automaton with weights in $S$ computing $f$) and hence by assumption, strongly guessable over $S$. Suppose for contradiction that $f$ is not strongly guessable over $S^{\prime}$. Then there exists a finite set $Q$ such that for all finite sets $T$ there exists $\underline{\lambda}_{T}\in\Lambda_{Q,T}\setminus\Lambda_{Q}$, where these solution sets are considered over $S^{\prime}$. However, noting that the coefficients $\underline{\lambda}_{T}$ lie in $S$, this directly contradicts that $f$ is strongly guessable over $S$. The final statement follows from Remark 3.10 together with the fact that any integral domain embeds into a field (its field of fractions). ∎
Remark 4.6.
Suppose $S^{\prime}$ is a subsemiring of $S$ and $f:\Sigma^{*}\rightarrow S^{\prime}$. Then we may also regard $f$ as a function with co-domain $S$, and as observed in the previous proof, if $f:\Sigma^{*}\rightarrow S^{\prime}$ is weakly guessable then $f$ is also weakly guessable when regarded as a function with co-domain $S$. However, if $f$ is guessable (respectively, strongly guessable) when considered as function with co-domain $S^{\prime}$, it need not be guessable (respectively, strongly guessable) when considered as a function with co-domain $S$. For example, the function computed by the automaton depicted in Figure 1(b) is strongly guessable when the weights (and hence the co-domain of the function) are considered to lie in the semiring $S^{\prime}=\mathbb{N}_{\rm max}$, but it is not guessable when the weights are considered to lie in $S=\mathbb{Z}_{\rm max}$ (see the functions denoted $f_{2}$ and $f_{3}$ in Appendix B for details).
We now apply the results and discussion of this section to determine whether Property A or B holds in the semirings used in our examples. The table in Figure 5 demonstrates that there are interesting and important semirings where Property B holds (i.e. all weakly (co-)guessable functions are strongly (co-)guessable) whilst Property A does not hold (i.e. there exist functions that are not weakly guessable). Since over such semirings the property of being strongly guessable is characterised in terms of automata (by Theorem 3.6), it is possible for the ‘teacher’ to select a strongly guessable target function to begin with, and in light of the discussion in Remark 3.26, there is then an algorithm (similar to that of [36], but using a different update strategy) that terminates with success for all such target functions with a closure strategy.
A B
$\mathbb{B}$ Yes Yes
$\mathbb{R}$ Yes Yes
$\mathbb{Z}$ Yes Yes
$\mathbb{N}$ No ($f^{\prime}_{1}$) Yes
$\mathbb{R}_{\geq 0}$ No ($f^{\prime}_{1}$) Yes
$\mathcal{P}_{\rm fin}(\Sigma^{*})$ No ($f^{\prime\prime}_{1}$) Yes
$\mathbb{N}_{\max}$ No ($f_{1}$) Yes
$\mathbb{Z}_{\max}$ No ($f_{1}$) No ($f_{2}$)
$\mathbb{R}_{\max}$ No ($f_{1}$) No ($f_{2}$)